OpenStudy (anonymous):

Which conjunction or disjunction is equivalent to the open sentence |a – 8| ≤ 3. A. a – 8 ≤ 3 and a – 8 ≥ –3 B. a – 8 ≥ 3 and a – 8 ≤ –3 C. a – 8 ≤ 3 or a – 8 ≥ –3 D. a – 8 ≥ 3 or a – 8 ≤ –3

OpenStudy (tkhunny):

It's "and". That eliminates two choices. Why would it be <3 with the absolute values but >3 without them? This eliminates the rest of the confusion.

OpenStudy (anonymous):

i dont really get it

OpenStudy (tkhunny):

\(|a-8|\le 3\) This is equivalent to \(-3 \le a-8 \le 3\)

OpenStudy (anonymous):

ok..but i dont see that in the answers?

OpenStudy (tkhunny):

That's because it is written as a conjunction, rather than a compound inequality. You should find the equivalent expression \(-3 \le a-8\; AND\;a-8 \le 3\)

OpenStudy (anonymous):

can you try explaining how to find the answer for questions like this?

OpenStudy (tkhunny):

It easy enough with ONE concept. |something| < is a Conjunction -- Two regions that overlap and we want the intersection. |something| > is a Disjunction -- Two regions that do not overlap and we want both pieces.